OpenStudy (anonymous):

Say i got a slope of -2. But the line is negative, would those two negatives cancel out to make a positive 2?

OpenStudy (anonymous):

Will medal and fan :) :)

pooja195 (pooja195):

so something like -2/-1?

OpenStudy (anonymous):

here hang on

OpenStudy (anonymous):

what is the slope of the linear function? i got -2. The points on the line are (0,2) and (4,0). It says that the line has a negative slope, but it also says my answer is wrong

OpenStudy (anonymous):

sorry i can't copy/paste the actual question it wont let me

OpenStudy (anonymous):

wait no the slope is -1/2

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{0-2}{0-4}=slope\] \[\LARGE \frac{0-2}{0-4}=-2/-4\] -2/-4=1/2

OpenStudy (anonymous):

ohhhh okay thanks :)

pooja195 (pooja195):

Does it make sense?

OpenStudy (anonymous):

yes it does :) it just clicked in my brain

pooja195 (pooja195):

xD good! :) and yw

OpenStudy (anonymous):

thanks so so much :)

pooja195 (pooja195):

No problem :)